(2x – 6)2 i have to find the square
Found 2 solutions by user_dude2008, stanbon:
Answer by user_dude2008(1862)

(2x - 6)^2 = (2x - 6)(2x - 6)=4x^2-12x-12x+36 = 4x^2-24x+36

Answer: (2x - 6)^2 = 4x^2-24x+36

Answer by stanbon(75887)

(2x – 6)^2
---
= (2x-6)(2x-6)
Use FOIL:
F: 4x^2
O: -12x
I: -12x
L: 36
===============
Combine to get:
= 4x^2 - 24x + 36
=======================
Cheers,
Stan H.